Hello i have been trying a few ways to try and do this sum but i dont seem to get it right.I tried it 4 different ways and im not getting a sensible answer.
this is what i did for solving linear quadratic equations but stopped because it doesnot look like im doing it right:
x^2=2y+10
3x-y=9
Solve eq 2 for y
3x-y=9
-y=9-3x
subs -y=9-3x
x^2+3x=2-9
x^2+3x=-7

Answer by John10(297)

You have the great idea by solving for y in the second equation but you are unfinished when you solve for y from this step:
-y = 9 - 3x
You have to divide both sides by -1 to solve for y:
y = 3x - 9
Now you will substitute y = 3x - 9 into the first equation:
x^2 = 2(3x - 9)+ 10
x^2 = 6x - 18 + 10
x^2 - 6x + 8 = 0
(x -4)(x-2) = 0
x = 2 or x = 4
Now we will substitute those two values to find two values of y:
If x =2 then y = 3(2) - 9 = 6 - 9 = -3
If x = 4 then y = 3(4) - 9 = 12 - 9 = 3
So we have two solutions: (2, -3) or (4, 3)
Hope it helps:)
John10